When using the addition method, I don't understand how to solve this problem:
4x+5/2y=11
3x-4/3y= -11
I did what my textbook suggested and tried to eliminate fractions, but ended up with something even more difficult:
24x+15y=11
18x-8y= -11
Please help I'm not sure what I should do!
Christina

Answer by mananth(16946)

(4x+5)/2y=11
multipluy by 2y
4x+5 = 22y
4x-22y=-5.....................(1)
(3x-4)/3y= -11
multiply by 3y
3x-4 = -33y
3x+33y=4.......................(2)
4 x -22 y = -5 .............1
3 x + 33 y = 4 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 2
12 x -66 y = -15
6 x + 66 y = 8
Add the two equations
18 x = -7
/ 18
x = - 7/18
plug value of x in (1)
4 x -22 y = -5
-14/9 -22 y = -5
-22 y = -5 -14/ 9
-22 y = -59/ 9
y = 14/47